An arithmetic sequence whose first term is 2 has the property that its second, third, and seventh terms are consecutive terms of a geometric sequence. Determine all possible second terms of the arithmetic sequence
1
Expert's answer
2020-12-15T02:11:19-0500

To find an arithmetic sequence is to find a and d.

Its second, third and seventh terms are 2+d, 2+2d and 2+6d . Therefore


(2+2d)2 = (2+d)(2+6d)


⇒ 4+8d+4d2=4+14d+6d2


⇒ 2d2+6d=0


⇒ d=0,−3


For Second terms d = 0 , -3


Second term 2 + d

i) 2 + 0 = 2

ii) 2 + (-3) = 2-3 = -1


2 , -1 are the possible second terms